### v2

# Think Outside the Data: Colonial Biases and Systemic Issues in Automated Moderation Pipelines for Low-Resource Languages

Abstract
Most social media users come from non-English speaking countries in the Global South, where much of harmful content appears in local languages. Yet, current AI-driven moderation systems struggle with low-resource languages spoken in these regions. This work examines the systemic challenges in building automated moderation tools for these languages. We conducted semi-structured interviews with 22 AI experts working on detecting harmful content in four low-resource languages: Tamil (South Asia), Swahili (East Africa), Maghrebi Arabic (North Africa), and Quechua (South America). Our findings show that beyond the well-known data scarcity in local languages, technical issues—such as outdated machine translation systems, sentiment and toxicity models grounded in Western values, and unreliable language detection technologies—undermine moderation efforts. Even with more data, current language models and preprocessing pipelines—primarily designed for English—struggle with the morphological richness, linguistic complexity, and code-mixing. As a result, automated moderation in Tamil, Swahili, Arabic, and Quechua remains fraught with inaccuracies and blind spots. Based on our findings, we argue that these limitations are not just technical gaps but reflect deeper structural inequities that continue to reproduce historical power imbalances. We conclude by discussing multi-stakeholder approaches to improve automated moderation for low-resource languages.
1 Introduction
The largest and fastest-growing user bases of social media companies come from the Global South, where billions of users generate content in their local languages. This growth has fueled a surge in non-English harmful content—such as misinformation, hate speech, and incitement to violence—contributing to severe human rights violations across the region (Samuels 2020; Milmo 2021; Yibeltal and Muia 2023). Yet, tech companies tend to prioritize moderation for English-speaking users in the West (Legon and Alsalman 2020; Popli 2021), leaving harmful content in languages spoken in the Global South largely unchecked. This neglect has deepened social harms (Nigatu and Raji 2024) and political divides across the Global South (Samuels 2020; Milmo 2021). Simultaneously, flawed moderation systems often misclassify and remove benign content in these languages, silencing marginalized voices and restricting freedom of expression (Elswah 2024b).
Historically, many languages spoken in the Global South are considered “low-resourced” due to the lack of high-quality datasets needed for training AI models (Rowe 2022; Nicholas and Bhatia 2023; Nigatu et al. 2024), which serve as the backbone of moderation infrastructure. However, data scarcity tells only part of the story. Economic and political oppression, insufficient human expertise, and limited access to digital infrastructures further exacerbate the “low-resourcedness” of these languages (Nigatu et al. 2024). Moreover, framing the problem solely as one of data scarcity overlooks broader challenges across the moderation pipeline, such as annotation, model training, and deployment. To address this critical gap, we examine the systemic barriers hindering equitable moderation for low-resource languages and explore actionable pathways to improve these systems. Specifically, we ask:
  • RQ1:
    What systemic barriers impact automated moderation pipelines for low-resource languages?
  • RQ2:
    How might we improve automated moderation for low-resource languages?
To address these questions, we conducted semi-structured interviews with 22 AI researchers and practitioners, specializing in harmful content detection and developing automated tools for diverse low-resource languages that have poor moderation support. These are: Tamil from South Asia, Swahili in East Africa, Maghrebi Arabic from North Africa, and Quechua in South America.
Our findings reveal a spectrum of systemic issues beyond data scarcity impacting the automated moderation pipeline for low-resource languages. Many participants criticized tech companies’ data restriction policy for hindering moderation research in the Global South. They pointed out that company’s use of biased machine translation systems, Western-centric toxicity models, and poor language detection tools—overlook the cultural nuances of online harms and language evolution in the Global South. They emphasized that even with more data, current English-centric design of preprocessing techniques (e.g., tokenization, stemming) and language models disregard the linguistic diversity, morphological complexity, and dynamic evolution of languages through code-mixing and code-switching, which are often absent in English. For instance, unlike English which has a relatively fixed word order (Bender 2009), Tamil, Swahili, Arabic, and Quechua have agglutinative property, meaning they can form thousands of complex words from a single root. Data-driven models primarily trained on English typically fail to infer these linguistic properties that do not exist in English. As a result, words that frequently appear in sexual harassment, such as Tamil word Mualichhu (meaning, n**ples) incorrectly gets stemmed to Mulai- (meaning, sprout) and goes undetected by models.
Drawing on these findings, we use coloniality as a lens to critically examine how tech companies perpetuate digital colonialism (Kwet 2019), prioritizing profit over user safety in less profitable markets in the Global South (Nicholas and Bhatia 2023). These companies not only monopolize the data extracted from next billion users in the Global South (Coleman 2018; Couldry and Mejias 2019) but also rely on biased data sources for moderation—reinforcing harmful narratives about formerly colonized populations. We highlight how current English-centric design of one-size-fits-all moderation tools reinforces colonial impulse by ignoring the linguistic diversity of Global South languages. We argue that improving moderation for these languages requires more than technical fixes, as competing stakeholder priorities demand deeper systemic changes. The key contributions of our work are as follows:
  • A qualitative study that uses coloniality as a lens to provide a critical and nuanced understanding of how historical power imbalances disproportionately affect automated moderation pipelines for diverse low-resource languages in the Global South.
  • An outline of paths forward, acknowledging the complexity, practical constraints, and systemic issues to improve moderation for low-resource languages.
2 Related Work
In this paper, we situate our work first by discussing existing content moderation literature focusing on the Global South. We then describe scholarly work investigating colonial biases in content moderation systems.
2.1 Content Moderation in the Global South
Content moderation refers to reviewing user-generated content to see if it aligns with tech company’s policies on what content should or should not be allowed on their platforms. Most tech companies moderate content using a combination of manual human reviews and automated AI models (Gorwa 2019). However, these companies often lack financial incentives to invest in moderation resources for less profitable markets in the Global South (De Gregorio and Stremlau 2023; Nicholas and Bhatia 2023). For instance, Meta funnels 87% of its global misinformation budget to the United States (US), despite Americans comprising only 10% of its user base (Popli 2021). The disparity is even more glaring when tech companies swiftly respond to harmful content from European countries that either offer strong economic incentives (De Gregorio and Stremlau 2023) or are of geopolitical interest to the US (e.g., Russia-Ukraine war) (Meta 2022). In contrast, tech companies have been less proactive in countering disinformation campaigns and extreme speech festering in many Global South countries (Milmo 2021; Wong and Ernst 2021; Wong and Harding 2021; Yibeltal and Muia 2023), while unjustly removing culturally appropriate and politically legitimate content from this region (Elswah 2024b; Shahid and Vashistha 2023).
The inability of tech companies to accurately and fairly moderate content in the Global South is often attributed to their reliance on automated moderation systems, trained on data-rich languages like English and a handful of European languages (Nicholas and Bhatia 2023; De Gregorio and Stremlau 2023). Prior research highlights that the lack of data in low-resource languages hinders the development of robust NLP technologies for detecting harmful content in these languages (Nicholas and Bhatia 2023; Nigatu et al. 2024). In contrast, little attention is given to other critical stages of automated moderation pipelines, such as who annotates what is harmful or what assumptions are made about deploying these models in complex, low-resource environments. To address this critical gap, we examine the systemic challenges AI researchers and practitioners encounter at various stages of automated moderation pipelines when developing moderation technologies for low-resource languages in the Global South. We now present scholarly work critically examining systemic issues in content moderation systems through the lenses of power and control.
2.2 Coloniality in Content Moderation
Coloniality perpetuates historical power imbalances through extraction, enslavement, and appropriation (Mbembe 2016). Decolonial scholars argue that the colonial structures persist today by exploiting the resources and labor of historically colonized populations while reinforcing Western dominance in governance and knowledge production (Quijano 2000, 2007a). Thus, decolonial computing critically examines who participates in computing, where it occurs, and how it shapes both knowledge (epistemology) and existence (ontology) from the perspective of those at the margins of the modern world system (Ali 2016). Whereas, postcolonial scholars conceptualize coloniality in computing as when technologies designed in the West, with Western values encounter diverse cultures (Irani et al. 2010).
Several scholars have critically examined content moderation systems through the lens of coloniality. For instance, Shahid and Vashistha (2023) use decoloniality as a lens to highlight how tech companies frequently impose Western values as global community standards, disregarding local socio-cultural norms when assessing online harms in the Global South. They draw parallels between Western-centrism in community guidelines and the way colonial powers systematically suppressed Indigenous and marginalized communities’ diverse ways of being, while imposing Euro-modern rationality as the only legitimate way (Said 2000; Gramsci 2020; Quijano 2007a). Similarly, Siapera (2022) provides a decolonial critique of how tech companies dismiss input from racialized users when shaping policies on racist hate speech. She argues that the company’s race-blind moderation policies mirror colonial legacies, where the identities and lived experiences of racialized communities are considered inferior (Quijano 2007b) and criminalized through “neutral” technologies (Benjamin 2023).
In addition, scholars draw attention to how tech companies build AI models for moderation on the backs of low-wage moderators and marginalized communities, whose labor and trauma fuel training datasets (Siapera 2022; Shahid and Vashistha 2023; Elswah 2024a). They point out that Western tech companies treat moderators from the Global South as dispensable, often concealing the true nature of the work during recruitment, and avoiding liability for the harms these moderators experience (Ahmad and Krzywdzinski 2022; Elswah 2024a).
Moreover, errors in moderation systems disproportionately affect marginalized communities, whose voices have been historically silenced. For example, AI models have been shown to drive systemic racism and heteronormative patriarchy by erroneously labeling Black and queer vernacular as toxic (Bhattacharyya 2018; Sap et al. 2019; Mohamed, Png, and Isaac 2020). Current moderation systems, shaped by a Western perspective, frequently misclassify innocuous and culturally appropriate content in non-Western contexts as harmful, while failing to detect actual harmful content (Shahid and Vashistha 2023). For example, Google’s Perspective API underestimates the toxicity of extreme speech in Swahili and Hindi, but rates similar content in Western languages, such as English and German more accurately (Udupa, Maronikolakis, and Wisiorek 2023). Udupa, Maronikolakis, and Wisiorek (2023) stressed that these errors persist because current moderation systems inherit Eurocentric colonial frameworks that rationalize uneven allocation of corporate resources for content moderation across different geographies and language communities.
We contribute to this growing body of work by critically examining the often-overlooked socio-political issues embedded at different stages of automated moderation pipelines, spanning data collection, labeling, cleaning, model training, and evaluation. As most studies overlook socio-political contexts when discussing power asymmetry within the AI pipeline (Ovalle et al. 2023)—to address this gap, we interrogate the normative assumptions and design paradigms underpinning AI-driven moderation technologies through the lenses of power and control. Drawing on interviews with AI experts, we explore how prevailing language technologies shape automated moderation of online harms in diverse Global South contexts (RQ1) and how moderation practices might be improved for these languages (RQ2).
3 Methodology
To examine disparities in automated moderation pipelines, we interviewed 22 AI researchers and practitioners specializing in automatic detection of harmful content in diverse low-resource languages spoken across the Global South.
Low-Resource Languages. We selected four linguistically diverse languages from different parts of the Global South. These are: Tamil from South Asia, Swahili from East Africa, Maghrebi Arabic from North Africa, and Quechua from South America (see Table 1 in Appendix). All these languages are considered low-resourced, despite being spoken by millions of people. UNESCO even declared Quechua as a vulnerable language due to systemic discrimination against Indigenous Quechua speakers in South America (Bank 2014). Due to limited resources, moderation errors are typically high for these languages. For instance, tech companies have repeatedly failed to address ethnic hate speech in Swahili (Witness 2022) and harmful content in Arabic (Elswah 2024b), while unjustly removing Tamil news articles as dangerous speech (Biddle 2022) and shadowbanning Arabic content on Palestine (Elswah 2024b).
Participants. We recruited people, who either (1) worked on automatic detection of harmful content, or (2) developed language models and tools in Tamil, Swahili, Maghrebi Arabic, or Quechua. We used purposive and snowball sampling to recruit 22 participants. Among them, six specialized in Tamil, six in Swahili, five in Maghrebi Arabic, and three in Quechua. Most of them (n=15) were native speakers of one of these languages. Many of our participants were affiliated with academia (n=13) and trust and safety teams at Meta, OpenAI, and TikTok (n=4). Some worked for trust and safety vendors, who built moderation tools and datasets for different clients (n=3) and local AI startups (n=4). Some participants held multiple roles. Five self-identified as women and the rest as men. All participants had experience living in the Global South, such as Kenya, Tanzania, India, Sri Lanka, Peru, Morocco, and Egypt. Half of them were affiliated with Western institutions and were based in North America and Europe during the interview. See Table 2 in Appendix.
Data Collection and Analysis. We conducted semi-structured interviews with the participants via Zoom. Each interview lasted for around 40-60 minutes. The semi-structured interviews focused on the collection, annotation, and preprocessing of the data in low-resource languages as well as model development. We also asked in detail about the models and tools they used to detect harmful content, and the reliability and performance of those models and tools. The participants also reflected on biases and challenges they encounter throughout the process and discussed ways to address them. After each interview, we iteratively refined our interview protocol, stopping when the responses reached saturation. After obtaining ethical approvals from IRB, we conducted the interviews in English and audio recorded with the consent of participants. We offered a modest compensation to the participants with \$100 Visa gift cards.
We transcribed the interviews, performed iterative open coding following reflexive thematic analysis (Braun and Clarke 2006), and continuously refined the emerging themes. Our coding process resulted in 441 codes, iteratively merged into 23 subthemes (e.g., model performance, annotation challenges)–which we mapped into different stages of automated moderation pipelines.
4 Findings
Figure 1: Issues affecting different stages of automated moderation pipeline for low-resource languages.
In this section, we outline systemic issues in moderating content in low-resource languages throughout automated moderation pipeline, spanning data curation (4.1), annotation (4.2), preprocessing (4.3), and model training (4.4).
4.1 Barriers to Access Datasets on Harmful Content
To detect harmful content, most participants needed large volumes of labeled data to train AI models, for which they often relied on user-generated content on social media platforms, such as Facebook, X, YouTube, and Reddit. Industry practitioners shared that their research and product teams have easy access to user-generated content on their platforms. In contrast, academic researchers pointed to structural barriers in studying emerging trends in online harms due to the lack of public datasets in low-resource languages and restricted access to social media data. For example, in 2018–19 when Twitter allowed free API access, it restricted researchers from accessing data older than two weeks. P14, an academic researcher working in Swahili commented:
People frequently used the word ‘madoadoa’ [spots] to spew hatred and violence during the 2007-08 Kenyan election. But that changed in the 2022 election. Bad actors appropriated the popular song ‘sipangwi’ [I am not told what to do] and its plural form ‘hatupagwingwi’ to spread hatred. Unfortunately we neither have access to recent nor past data to study how hate speech tactics have evolved.
To bypass API restrictions, many researchers and small trust and safety vendors used open source scrapers to collect user-generated content. However, they noted that these scrapers struggle to capture romanized and code-mixed content in their target languages, frequently misidentifying it as English because of Latin alphabets used in writing.
Recently tech companies, such as Meta, X, and Reddit have blocked these scrapers and API access to user-generated content, such as through Meta’s CrowdTangle program (Mehta 2023; Bellan 2024; Stokel-Walker 2024; Perez 2024). These restrictions along with the lack of available datasets significantly reduced the ability of trust and safety vendors and academic researchers to study disinformation and hate speech campaigns in at-risk countries in the Global South. P18, a trust and safety practitioner from a major social media company remarked:
After ChatGPT came out, companies are cautious of publicly sharing their data given the competition to develop their own language models. That’s why we no longer see that openness around sharing data.
Given the difficulties in accessing and curating datasets, some researchers ceased studying online harms in their regions. Others turned to alternative sources, such as using datasets from shared tasks at NLP conferences, manually collecting online posts, surveying local communities to collect harmful content, or relying on voluntary data donations from WhatsApp groups. However, these methods proved time-consuming and produced small, sporadic datasets, often inadequate for training AI models effectively.
Some participants argued that tech companies must grant researchers access to user-generated content. While they recognized the privacy and ethical concerns associated with data sharing practices, they demanded equitable access to these data because some platforms like TikTok only provide API access to researchers in the US and Europe (TikTok 2024). Given these systemic discrepancies in company’s data sharing practices, African researchers started creating and joining grassroots efforts, such as Masakhane and Tanzanian AI, to establish ownership of the data generated by users in their communities. Academic researchers criticized tech companies for mishandling harmful content in their region despite controlling user data. P19, who worked at a major tech company commented:
When I worked at [redacted], the trust and safety team prioritized the US. These for-profit corporations derive most of their revenues from markets that are outside the Global South. Although Europe has strong regulatory policies, those markets are important to the company. So the prioritization simply reflects that.
Participants highlighted several issues in how tech companies address data scarcity in low-resource languages. They shared that the keyword based filtering commonly used by tech companies to identify harmful content often falls short, as it ignores dialectical variations and treats these languages as monolithic. They also criticized the use of machine-translated texts as a workaround for limited data due to biases in tech company’s machine translation tools. For instance, Kenyan Swahili researchers observed that Google Translate frequently incorporates outdated Sheng—a cerole blending Swahili and English—but fails to support its modern variants like Shembeteng. In contrast, Tanzanian researchers criticized Google Translate for favoring Kenyan Sheng, overlooking the purer Swahili spoken in Tanzania. Similarly, Quechua researchers highlighted tech company’s problematic reliance on outdated sources, such as Bible translations and the diaries of colonial-era priests, to compensate for the lack of digitized content in Indigenous languages like Quechua. An industry practitioner also noted that their company relied on old Arabic dictionary due to limited datasets in Maghrebi Arabic.
Our study participants stressed that the models relying on outdated corpora and biased machine translations are ill-equipped to address the evolving nature of hate speech online. Additionally, a trust and safety practitioner highlighted logistical and legal barriers that often impede moderation efforts within the company. They noted that while certain open-source multilingual model achieved better machine translation in low-resource languages, the company could not deploy it to improve moderation due to licensing issues related to the model’s training data.
Small AI startups and trust and safety vendors shared that big tech companies often showed interests in their datasets and tools developed for low-resource languages, but only if they worked for free. Many researchers demanded tech companies to invest in low-resource languages and strengthen grassroots, local research capacity to address online harms in the Global South. P9, a Quechua linguist expressed:
They [companies] should work with us, indigenous Quechua people, to build corpuses instead of taking the shortcut by using machine-translated texts. We found rule-based translation that incorporates grammatical knowledge works better for Quechua than stochastic methods, which require lots of data that do not exist in Quechua. When we contacted Google, they proposed us to work voluntarily. So, I’m worried they will try to appropriate our free labor.
These findings show that tech companies’ reluctance to invest in data sources in the Global South and gatekeeping of user-generated content on their platforms amplify prevailing data scarcity in low-resource languages, disrupting grassroots efforts to detect online harms in the Global South.
4.2 Difficulties in Annotating Harmful Content
Annotation involves labeling the data to train AI models to predict whether a content is harmful, and if so, identify the specific type. Tech companies frequently rely on manual reviews done by human moderators to train their AI models. Participants who worked at US-based tech companies shared that their companies partnered with vendors in the Global South to annotate large scale user-generated content. One of the participants shared that their company often assigned Kenyan moderators to annotate different dialects of Swahili, even when the moderators didn’t know those dialects. Company’s efforts to assign content to appropriate moderators with language expertise often fail because language identification technologies perform poorly in low-resource languages. Participants further stressed that companies have always underfunded annotation efforts for languages spoken in “less profitable regions.” P19, who worked at a social media company shared:
During Arab Spring, [redacted] had only two Arabic speaking moderators. There’s so much diversity in the Arab world– it’s unlikely that the two moderators will get the full context of Arab Spring in Tunisia or Green Movement in Iran. Although a lot has changed since then, the core structure and issues remain the same.
As a result, participants observed that tech companies lack a deep understanding of ground realities, social and cultural norms, and linguistic nuances of low-resource languages, which significantly hinders company’s ability to address harmful content in the Global South. They stressed that tech companies should “give the Global South a seat at the table” when defining hate speech. P12, an academic researcher working on Swahili remarked:
It matters who is defining hate speech. We noticed that people use ‘US’ vs. ‘Them’ narrative to spread ethnic hate speech and superlatives to express supremacist views. We developed our annotation framework to capture these cases. Since Twitter did not remove these tweets, their definition of hate speech must be different. By allowing these posts Twitter is reinforcing stereotypes about Africans being violent.
To ensure that the annotated datasets capture local sensitivities, local researchers often involved linguists, activists, and affected communities to inform their annotation guidelines. P3, a researcher working on Tamil explained:
It’s important to consider intersectionality when annotating hate speech in multicultural environments like India, where caste, religion, and gender are intertwined. For example, we found “shuttlecock” [badminton cork] is used as a derogatory term against Muslim women who wear burka. Our team of feminist activists, experts on gender studies, and survivors of harassment helped us annotate coded hate speech that are both misogynist and Islamophobic. Similarly, there were innocuous comments like “you are my sweetheart.” When companies recruit gigworkers who are usually male, they would rate this as harmless. But since we worked with victims of sexual harassment and recipients of such comments, they could recognize these messages are part of broader harassment Indian women face online.
While researchers appreciated the value of involving community partners in data annotation, they also struggled with requisite funding to sustain these partnerships, provide annotator training, and maintain the quality of annotation. The lack of funding also forced them to rely on undergraduates to annotate hate speech and toxic datasets, without being able to provide mental health support for these students. P11, an academic researcher working on Quechua shared:
Very often the dataset we are creating is the first of its kind in Quechua. Although experts and community members are willing to help voluntarily, it’s difficult to sustain their free labor in the long run to annotate large volumes of data. So, we often strategize to annotate only a subset of data. We can’t rush people to annotate faster because they are helping out of generosity. Thus, it takes months to annotate anything.
To make the most of limited annotation resources, researchers often used sentiment and toxicity analysis tools to find negative content, reducing the sample size for manual annotation. However, they noted that existing free and proprietary tools from tech companies often lack cultural nuances. P13, a researcher specializing in Swahili elaborated:
In America, people casually use the word “dawg” to refer to buddy but in Kenya calling someone dawg will be disrespectful. Similarly, in America people think calling “fat” is body shaming. In Africa fat is considered beautiful and opulent. But Google’s perspective API missed these cases by applying American scale.
Tamil researchers shared that they lose valuable annotation time and budget when manually verifying target languages in scraped corpuses. Existing language identification tools have poor coverage for most low-resource languages. Thus, these tools often fail to separate code-mixed Tanglish (Tamil-English) from Kannada-English or Telugu-English because Tamil, Kannada, and Telugu often share words with same roots. Similarly, Maghrebi Arabic researchers reported that these tools often fail to differentiate between Arabic, Farsi, and Urdu due to overlap among their scripts.
These findings show that poor coverage of low-resource, non-English languages in current AI advances complicates the annotation process for these languages. Moreover, grassroots annotation efforts are limited due to chronic underfunding. Furthermore, despite having ample resources, tech companies often fail to capture the cultural nuances of online harms due to inadequate engagement with stakeholders and affected communities in the Global South.
4.3 Preprocessing Challenges for Harmful Content Detection
Preprocessing involves cleaning and transforming raw data in a suitable format to train AI models. Our participants faced several challenges when applying existing preprocessing techniques on low-resource languages.
Tokenization. Tokenization is a crucial preprocessing step where the text is segmented into smaller units, such as words or subwords, to enable models to process and analyze language effectively. Several participants shared that the multilingual AI models they used for detecting harmful content, such as BERT and RoBERTa use frequency-based tokenization algorithms, such as WordPiece and BPE. These algorithms generate tokens based on the frequency of words or co-occurring character pairs in the dataset. However, participants noticed that this technique performs poorly on Tamil, Swahili, Maghrebi Arabic, and Quechua texts because these languages have richer and more complex morphology than English. They explained that Tamil, Swahili, Arabic, and Quechua have agglutinative properties, forming complex words by combining multiple morphemes (i.e., the smallest unit of meaning), with each morpheme retaining its original meaning. For example, the Quechua word ‘rimanqakuma’ (meaning, they will definitely speak) consists of three morphemes: ‘rima-’ (meaning, to speak), ‘-nqa’ (refers to future tense) and ‘-kuma’ (signifies emphasis). The final meaning is directly derived from these constituent morphemes. P9 elaborated further stressing the need to derive morphemes correctly during tokenization:
Frequency based tokenizers have been designed considering English as a model language. Since English is data-rich, frequency based method really works well. But for low-resource, agglutinative languages it creates illegible tokens by wrongly splitting the morphemes. If we train models with wrongly split tokens, the models won’t derive correct embeddings. Instead, when we used linguistically motivated tokenizer, the performance significantly improved for Quechua in downstream tasks.
Maghrebi Arabic NLP researchers also noted that using specialized morphological and monolingual tokenizers improve sentiment analyses for diverse low-resource languages, typically underrepresented in multilingual models. Swahili researchers further highlighted the challenges of tokenizing code-mixed hashtags that are often used to incite attacks while evading detection by platforms. For example, in the Sheng hashtag #TupataneTuesday (meaning, let’s meet each other on Tuesday), used by protesters, the Swahili word Tupatane must be correctly segmented into its morphemes: Tu- (we), -pat (to meet), -ane (each other). However, poor performance of language identification technologies on code-mixed texts complicates the selective application of tokenization algorithm based on language.
Normalization. Researchers also identified challenges in the normalization process performed by tokenizers, where words are converted to their standard forms before tokenizing (e.g., baaaad is normalized to bad). Some participants reported that non-standard spelling of agglutinative words causes confusion during normalization. P6, a Tamil researcher from Sri Lanka explained:
In Tamil, ‘Amma’ means Mother and ‘Ama’ means Yes. On social media people often enthusiastically write Ama as ‘Aammaa’ (similar to Yeessss) or distort the word Amma as ‘Aammaa’ in gendered slurs. The model often makes errors while normalizing such cases and fails to flag offensive language.
Stemming and Lemmatization. These steps are performed to reduce words to their meaningful roots before training models (e.g., beautiful and beautify are reduced to beauty). Several participants reported facing challenges because existing tools have higher error rates in complex agglutinative languages, where “each root can take thousands of inflected forms”, than morphologically simpler languages like English. P6 further described:
Both understemming and overstemming of complex Tamil grammar can cause error in detecting offensive language. Words like Mulaicchu (meaning, n**ples) often wrongly gets stemmed to Mulai- (meaning, sprout) and then gets ignored by model.
Parts-of-Speech Tagging. Some participants reported that since most models are trained on English, which is a subject-verb-object (SVO) language, it leads to errors on languages that follow subject-object-verb (SOV) structure. Therefore, they performed parts-of-speech (POS) analysis during data preprocessing to give models additional contexts about derogatory adjectives and verbs aimed at individuals or groups (nouns). For example, in Nāyai seruppāla aṭikkaṇum (meaning, beat the dog with sandals) the object (noun) Nāyai appears before the verb aṭikkaṇum. However, researchers faced several challenges in detecting POS due to code-mixing. P2, an academic researcher shared:
When I started doing NLP research in early 2000, there was no POS tagger for Tamil. There was barely any dataset to work with. We built corpuses from scratch and worked with linguists to annotate complex Tamil vocabulary. But the POS tagger based on monolingual Tamil does not work well on Tanglish from social media. Although many frame code-mixed data as problematic and low-quality, this is the reality of how social media users from non-English speaking countries write online. Handling code-mixing is very challenging. But we don’t have access to code-mixed data from social media since they stopped access.
These findings show that current preprocessing techniques, predominantly developed with English in mind, do not account for the linguistic diversity of morphologically rich and code-mixed nature of languages in the Global South, reflecting historical imbalances in linguistic and technological priorities.
4.4 Challenges in Developing and Training AI Models for Harmful Content Detection
After preprocessing data into a standard format, it is fed into AI models for training and detecting harmful content. Participants reported using various multilingual language models, such as Google’s mBERT, Facebook’s XLM-RoBERTa, and AI4Bharat’s IndicBERT for detecting harmful content. However, they noted that these language models perform poorly on low-resource languages. They pointed out that although these data-driven models are designed to be language-agnostic, being primarily trained on high-resource languages like English, they better learn the simpler morphology and fixed word orders of English. In contrast, data sparsity in low-resource languages limits these models’ ability to fully capture the rich inflectional morphology, agglutinative property, complex grammar, and diverse word orders in languages that are linguistically distinct from English. P4, specializing in Tamil described:
English and Tamil are from different language families and Tamil has richer morphology than English. How can these models derive correct embeddings of complex Tamil words by computing from the point of view of English? That’s why IndicBERT doesn’t perform well. There, Hindi and Marathi are from the same family but Tamil is a Dravidian language. So without considering the specifics of language families, you can’t get performance improvement.
Researchers cautioned that adding data from multiple languages can degrade model’s performance in both low-resource and high-resource languages due to limited model capacity, a phenomenon known as the “curse of multilinguality.” Additionally, they criticized how AI models cannot handle diversity in annotations, especially for languages like Tamil and Swahili that have tens of dialectical variations. P15, a startup founder focusing on Swahili AI explained:
In Swahili the word ‘right’ has at least 20 different transliterations depending on the context. Similarly, in my region, the word ‘Mathikkalla’ refers to ‘I could not recognize you’ but in other regions, the same word means ‘to neglect someone.’ So, annotators would label the same content differently depending on their region. This impacts offensive language detection because AI models flatten the diversity in annotation into a singular view.
Some researchers observed that large language models frequently misclassify code-mixed content during hate speech detection, especially when the spelling and words signal non-Western ethnicity. Trust and safety practitioners attributed these errors to a lack of diversity within tech companies and shared that very often their teams are linguistically and culturally homogeneous. They commented that company’s diversity efforts often end at recruitment; once hired, employees have to work following company’s priorities, which are typically centered around English. One practitioner from a US-based social media company remarked how this lack of diversity leads to biased models:
In Western media, Arabic phrases, such as “Allahu Akbar” [God is great] mostly appear in the context of terrorism. When companies train AI models on such articles, the models learn these negative associations. But there is none in these teams to inform that local people use these phrases to express everyday joy and sorrow, beyond the instances of extreme speech portrayed by Western media.
Industry practitioners shared that despite the shortcomings of large language models in low-resource languages, their companies are prioritizing AI models over alternative linguistic approaches they used in the past. They emphasized the advantages of using AI models for moderation, particularly in reducing the burden of tedious and distressing moderation work for humans. In contrast, AI researchers and practitioners working in the Global South highlighted their struggle in training billion parameter models due to a lack of funding, computational power, and appropriate hardware. For example, Swahili researchers and engineers shared that they could not buy GPUs in Kenya and Tanzania and had to rely on their contacts in the US to access these resources. Many pointed out that free resources from Google Colab and Kaggle are barely enough to experiment with, train, and deploy these language models. P12 commented:
We lack the necessary data, funding, and resources to build dedicated models for our languages. Our time is spent on scraping for little data and cleaning it. I hope we can decolonize NLP research on online harms, so that we no longer have to rely on technologies biased towards high-resource languages like English and developed for nations with lots of computing power.
These findings highlight that resource-intensive large language models, predominantly designed with an English-centric focus, are ill-equipped to address online harms in low-resource languages from the Global South, reflecting how the needs of these communities are usually sidelined in the development of AI-driven moderation technologies.
5 Discussion
While prior work attributes moderation challenges in low-resource languages to the lack of labeled datasets (Rowe 2022; Nicholas and Bhatia 2023), our study uncovers how socio-political factors in technology design exacerbate these issues. Our findings underscore how tech companies continue to rely on biased machine translation systems using outdated corpora instead of collaborating with experts and communities from the Global South—often appropriating their free labor. We reveal how these companies’ blanket data restrictions for building proprietary large language models aggravate data scarcity to address online harms in these regions (4.1). While prior studies report biased and opaque annotation practices among tech workers (Scheuerman and Brubaker 2024) and ML researchers (Geiger et al. 2020), we examine the structural factors enabling these issues. Socio-political issues, such as tech companies’ weak financial incentives to improve annotation for Global South languages and limited funding available to Global South researchers along with technical issues like Western-centrism in sentiment and toxicity models and treating code-mixed data as “poorer quality” when developing language detection tools—compromise the annotation processes (4.2).
Moreover, most studies explain away preprocessing and model building challenges in low-resource languages by highlighting data scarcity (Khan et al. 2023; Zhong et al. 2024). In contrast, our study questions the status quo that prioritizes data-intensive methods while overlooking alternative approaches that center linguistic diversity, morphological complexity, and dynamic evolution through code-mixing and code-switching—phenomena largely absent in English (4.3, 4.4). We provide concrete examples of how normative assumptions in technology design contribute to moderation errors in diverse Global South languages—that remain invisible when assessed solely through low accuracy rates. Our focus on diverse languages help us establish the systemic nature of moderation biases. In discussion, we probe deeper into these systemic inequities, unpacking their historical and socio-political roots—often overlooked in existing discourse (5.1). We then discuss approaches to improve moderation for low-resource languages while acknowledging the complexity of the issue (5.2).
5.1 Coloniality in Moderation Pipelines
Data Curation. Our data shows that tech companies lack interest to expend moderation resources for less profitable markets in the Global South. Our participants stressed that companies benefit by monopolizing user-generated data to train proprietary large language models, while restricting researchers’ access to the very data needed for detecting harmful content. For instance, shortly after Reddit locked public data (Perez 2024), it partnered with OpenAI to enable training ChatGPT on its content (OpenAI 2024). Similarly, Meta launched AI across Facebook, WhatsApp, and Instagram to train proprietary models on public posts without letting users opt out (Jiménez 2024), while simultaneously closing CrowdTangle that allowed researchers to access public content on Meta (Bellan 2024). Researchers criticized these blanket restrictions on public data as privacy washing, impeding trust and safety scholarship within academia and civil society (Arney 2024).
These restrictions disproportionately affect researchers and practitioners in the Global South, where datasets in non-English languages remain scarce. This data scarcity stems from colonial legacy that suppressed Indigenous and native languages in the Global South (Thiong’o 1986; Bank 2014; Obi-Young 2018; Kolli 2024) and deprioritized their digitization and technology development (Bird 2020; Schwartz 2022; Held et al. 2023; Ògúnrėm̀í, Nekoto, and Samuel 2023). The systemic omission affects all downstream NLP tasks in low-resource languages, including automated moderation– further hampered by data restriction imposed by tech companies.
Our participants highlighted that the data controlled by tech companies are generated through the unpaid labor of users in their communities. Coleman (2018) explains that Facebook introduced Free Basics initiatives in the Global South to extract data from the region’s next billion users, taking advantage of weak data protection laws and regulatory frameworks. Kwet (2019) likens this process to digital colonialism. He argues that much like colonizers who built railroads to extract material resources from colonies, tech companies control digital infrastructures in the Global South, reduce local communities to products rather than producers, and commodify their data for corporate profit.
Our analysis reveals that tech companies’ reliance on cheap web-scraped data, machine translations, and religious texts for low-resource languages (Kreutzer et al. 2022; Christodouloupoulos and Steedman 2015; Ghosh and Caliskan 2023)–introduces significant biases in moderation. This includes wrongly associating Arabic phrases with terrorism and normalizing extreme speech in African contexts. Such biases reflect digital orientalism (Alimardani and Elswah 2021), where colonial perspectives shape discriminatory narratives regarding the colonized ‘other’ (Said 1977). Likewise, the use of colonial-era texts to build Quechua datasets overlooks the historical role of colonial churches in suppressing Indigenous languages, while appropriating them only for cultural control (Heller and McElhinny 2017, p. 29). Thus, our findings highlight how colonial legacies continue to shape data sources used to study online harms in low-resource languages.
Annotation. Trust and safety practitioners in our study noted that tech companies lack economic incentives to recruit moderators and annotators with relevant expertise for content in the Global South. However, research shows that companies often outsource annotation tasks to the Global South for reviewing English-language content, exploiting low wages and weak labor protections  (Elizabeth Dwoskin and Cabato 2019; Elswah 2024a). This practice mirrors colonial exploitation, where the Global South workforce serves the interests of the Global North with little regard for local needs or equity (Posada 2021; Malik 2022).
Additionally, we found that limited funding in Global South institutions hinders grassroots efforts to annotate harmful content in local languages. Historically, resources extracted through colonial exploitation enabled Western nations to advance their scientific agenda and build extensive datasets (Schöpf 2020). Consequently, most misinformation research focuses on the West due to easy availability of annotated datasets in English. These systemic inequities, marked by resource scarcity in the Global South and tech companies’ disinterest in investing in these regions (Nicholas and Bhatia 2023; De Gregorio and Stremlau 2023)– further limit the availability of annotated datasets in low-resource languages.
NLP Tools Used in Moderation. Our findings underscore that current NLP technologies, primarily designed for English, overlook the cultural context, linguistic complexity, and evolution of languages in the Global South. For example, our participants reported that Google’s Perspective API misinterprets diverse notions of toxicity across different cultures. Similarly, Das et al. (2024) demonstrate that sentiment analysis tools for low-resource languages disproportionately associate negative sentiment with certain religious and national identities– replicating colonial hierarchies of division sowed by British rulers in the Indian subcontinent.
Decolonial scholars and historians have long documented the colonial project of standardizing European languages by creating dictionaries and grammars to assimilate Indigenous populations while suppressing local languages (Fishman 1989; Heller and McElhinny 2017; Anderson 2020; Fanon 2023). These forced affected communities to code-switch between native and European languages to navigate colonized spaces (Mufwene 2004). These legacies resulted in poor early support for non-Latin scripts online, continuing to hinder participation from speakers of many low-resource languages (van Esch et al. 2019; Held et al. 2023; Nigatu et al. 2024). This discrimination has forced non-English speakers to adopt romanization and code-mixing for communicating online (Held et al. 2023). However, the closed, proprietary language models, relying on sanitized datasets, disenfranchise local knowledge, impose Western normative values without empowering local communities to align the model to their own values, forestall alternative visions, and perpetuate colonial binaries that frame advanced technologies as rescuing “primitive” languages (Verran and Christie 2007; Bird 2020; Varshney 2024).
Primarily being trained on English, these language models perform well on languages that share important typological properties with English (Bender 2009; Arnett and Bergen 2024). Thus, these models fail to capture the elaborate morphology present in many low-resource languages. Historically, linguists considered agglutinative languages as “less evolved” than Western languages, such as Spanish, Greek, or German (Errington 2007). Bender (2009) critiques AI models for making assumptions about language structures that advantage some languages at the expense of others, highlighting their inherent lack of language independence. Scholars criticize such one-size-fits-all solutions for embodying “colonial impulse” that disregards the ecology of diverse languages and perpetuates colonial hierarchies (Dourish and Mainwaring 2012; Bird 2022). For languages spoken in the Global South, this translates to collapsing their linguistic diversity and complexity to a simplistic construct of data scarcity– often taken at the face value.
In sum, our findings show that existing challenges affecting automatic detection of harmful content in low-resource languages are often systemic and run deeper than the mere availability of data.
5.2 Considerations for A Path Forward
Tackling harmful content in low-resource languages is a complex issue shaped by conflicting interests and priorities across stakeholders. To begin with, private tech companies often consider it financially unviable to invest in moderation systems for low-resource languages even when these languages have millions of speakers (Nicholas and Bhatia 2023; De Gregorio and Stremlau 2023). Moreover, the ongoing deprioritization of trust and safety efforts within US tech companies undermines global accountability, prioritizing a US-centric vision of free speech (Scarcella 2024; Divon and Ong 2025). Academics also face disincentives. The time and effort required to create labeled datasets for low-resource languages (Sambasivan et al. 2021), combined with limited career payoffs and citation potential (Held et al. 2023), discourage research in this area. Governments in many Global South countries, frustrated by platforms’ failures to address hate speech and disinformation, often resort to censorship or criminalize political speech, further exacerbating the issue (De Gregorio and Stremlau 2023). Even when these governments mandate to store local user data within the country, they face pushback from US-based Silicon Valley lobbying (Kak 2020). On the other hand, civil society groups in the Global South frequently feel marginalized. Unlike their Western counterparts, they report limited influence, as tech companies often approach collaboration as a checkbox exercise rather than a genuine partnership (Centre 2024). Fully recognizing these issues as well as the constraints and complexities faced by all stakeholders, we outline some concrete steps to make content moderation more equitable.
Strengthening Local Research Capacity. Prior research highlights that when Global North institutions are funded to develop models for low-resource languages without involving local experts, they often fail in context-specific moderation tasks (Nicholas and Bhatia 2023). Bhabha (2011) argues that enhancing “national resources” of the Global South is essential to addressing the geo-politics of resource distribution and the transnational moral demands of redistributive justice. Therefore, governments, grant-making agencies, and research award programs by tech companies must invest in building self-sustaining, grassroots research ecosystems that actively engage local experts from the Global South. For example, the AI4D Africa program, funded by international governments and research institutes, supports the development of local AI research hubs and talent, empowering African researchers to lead projects that address their communities’ needs (IDRC 2024). Initiatives like Masakhane in Africa, AI4Bharat in India, and ARBML in the Arab World, which are democratizing AI research on low-resource languages, should be strengthened through targeted funding to amplify their impact.
Labeled Datasets. Social media companies should provide local researchers with access to de-identified data in low-resource languages. This would enable researchers to develop culturally and contextually appropriate labeled datasets and empower companies to address harmful content using these datasets. While companies frequently cite privacy concerns in data sharing, established practices from other fields suggest feasible solutions. For instance, the Yale Open Data Access (YODA) Project allows medical companies to securely share anonymized clinical trial data with vetted researchers for approved studies (Nicholas and Thakur 2022). Similarly, researchers recommend differential privacy techniques to protect personal information when sharing large datasets (Kapelke 2020; Garfinkel and Bowen 2022). Tech companies can adopt these strategies to enable secure and privacy-preserving access to data.
For languages with a significant digital presence, voluntary data donation by native speakers can be useful for grassroots researchers. For example, Garimella and Chauchard (2024) developed a data donation tool for closed WhatsApp groups while safeguarding the privacy of both donors and their contacts. In contrast, for Indigenous languages with limited digital presence, building respectful and equitable community relationships is essential, prioritizing local agency in community partnerships (Bird 2020). As participants highlighted the importance of diverse viewpoints when annotating intersectional hate speech, companies should actively seek out diverse annotators, prioritizing high recall to capture as many potentially harmful cases as possible (Parrish et al. 2024). While doing so, it is important to follow the best practices for supporting community labor when annotating harmful content by disclosing the task, offering opt-out options, providing well-being support, and monetary compensation  (Radiya-Dixit and Bogen 2024). For example, Karya—a nonprofit data company based in India—empowers disadvantaged communities through data annotation work and pays them nearly 20 times more than the local minimum wage (Perrigo 2023). Civil society groups should urge tech companies to recruit diverse moderators for local dialects, balance moderator’s workload when assigning traumatizing content, and ensure fare wages (Elswah 2024a).
Language-Aware Solutions. Since current NLP tools and language models are inherently English-centric, our participants recommended approaches that incorporate linguistic knowledge, such as using morphological segmenters instead of frequency-based tokenizers (Abdelali et al. 2021; Zevallos and Bel 2023), rule-based translations over stochastic machine translations (Sreelekha, Bhattacharyya, and Malathi 2018), and vector embeddings of local hateful phrases for detecting code-mixed hate speech (Devi, Kannimuthu, and Madasamy 2024). Some participants also discouraged using multilingual models for moderation given these models fail to infer correct linguistic knowledge for different language families. However, given the arms race among tech companies to develop large multilingual models (Gupta 2024), it is unlikely that they will shift to such linguistically informed solutions without regulatory pressures. Meanwhile, limited access to computing power in the Global South limits researchers’ capacity in training and experimenting with “language-specific” (monolingual) models and “language-aware” approaches that do not necessarily rely on vast datasets or huge computing power. Free computing resources provided by tech companies, such as Google’s Colab and TPU Research Cloud programs–remain inadequate for research purposes. Expanding access to these resources is critical to enable more equitable and inclusive moderation research in Global South languages.
Policy and Practice. For the data and tools created by local experts to have meaningful impact, they must be deployed to moderate local content. Cohere’s partnership with HausaNLP to integrate African language datasets into its multilingual Aya model (Radiya-Dixit and Bogen 2024) demonstrates the potential of such efforts. Similarly, the scales created by local researchers to evaluate model’s performance for detecting code-mixed hate speech (Das et al. 2022)–should be integrated into tech companies’ evaluation frameworks. Governments and civil society should create regulatory policies that require tech companies to prioritize local representation, data ownership, and community self-determination. These policies must articulate the specific harms caused by flawed content moderation systems in high-risk, under-resourced contexts, rather than simply applying Western frameworks of harm to regional trust and safety efforts (Kennedy and Campos 2025). Additionally, they should push for the inclusion of model performance metrics for low-resource languages in transparency reports, evaluated against locally defined benchmarks. Given the limitations of accuracy metrics in class-imbalanced content moderation contexts, regulatory frameworks should mandate the inclusion of more informative measures such as recall (% of correctly flagged harmful content) and precision (% of flagged content that is truly harmful) (Wei, Zufall, and Jia 2025). Such steps could surface the limitations of current models and incentivize progress toward better moderation of harmful content in underrepresented languages.
6 Research Positionality
All authors of this work come from historically colonized regions in the Global South and are native speakers of languages, which are considered “low-resourced.” All authors have extensive experience of doing critical research in diverse Global South contexts. Although none of the authors are affiliated with industry, our background in academia (computer science and communications) and civil society—enabled us to engage with participants both at technical and policy levels. Although we come from historically colonized countries, we are affiliated with academic institutions that have benefited from colonial expansion and were established using wealth derived from the forced appropriation of Indigenous lands. We acknowledge that these affiliations afford us research privileges—such as access to funding, institutional support, and global visibility—that are often inaccessible to many researchers based in the Global South. In this tension, we identify with Villenas (1996)’s notion of having “feet in both worlds,” as we simultaneously belong to communities shaped by colonial histories and to academic institutions that have profited from those same colonial legacies.
Similarly, we recognize that caste, religion, ethnicity, and other intersectional identity likely shaped the experiences of our participants although we did not explicitly collect such data. For instance, several Indian participants studying Tamil were likely from upper-caste backgrounds, some researchers studying Quechua were not of Indigenous origin, and many had greater access to resources due to Western affiliations compared to their counterparts in the Global South.
Following Fine (1994)’s self-reflexive approach, we acknowledge the intersections of privilege and marginalization among our participants. For instance, Quechua-speaking Indigenous researchers reported greater struggle in practicing their native language in academic spaces compared to Western researchers studying the same “low-resource language.” Similarly, the experiences of upper-caste AI experts would differ from those coming from lower-caste backgrounds–who might face additional challenges and lack social capital in pursuing AI research in Tamil. Swahili researchers and practitioners from Tanzania felt their Kenyan counterparts enjoyed more visibility since most tech companies have regional offices within Kenya. Additionally, although industry practitioners had greater access to computing resources, they reflected on their limitations within corporate financial infrastructures and lack of diversity within these organizations. As Haraway (2013) argues, knowledge production is always situated; the experiences of AI researchers and practitioners in our study form a “partial perspective”–shaped by their backgrounds and the struggle of researching marginalized, low-resource languages.
Appendix A Appendix
Table 1: Various characteristics of four low-resource languages featured in this study.
Tamil Swahili Maghrebi Arabic Quechua
Number of speakers 80 million 100 million 88 million 8 million
Geographic region
South Asia: Tamil Nadu
(India), Sri Lanka, etc.
East Africa: Kenya,
Tanzania, etc.
North Africa: Morocco,
Algeria, Tunisia, etc.
Andes: Bolivia,
Peru, Ecuador, etc.
Language family Dravidian Bantu Semitic
Quechuan
Grammar
Agglutinative, subject-
object-verb (SOV)
Agglutinative, subject-
verb-object (SVO)
Root based, verb-
subject-object (VSO)
Agglutinative, subject-
object-verb (SOV)
Colonial influence British
Portugese, German,
British, Arabic
French, Spanish, Italian Spanish
Table 2: Demographics of participants in our study.
Participant
ID
Language
Expertise
Role
Participant
ID
Language
Expertise
Role
P1 Tamil
Professor,
Startup founder
P12 Swahili Professor
P2 Tamil Professor P13 Swahili Master’s student
P3 Tamil PhD student P14 Swahili Professor
P4 Tamil Master’s student P15 Swahili Startup founder
P5 Tamil Software engineer P16 Swahili ML engineer
P6 Tamil Lecturer P17 Swahili Industry practitioner
P7 Indic languages Trust and safety vendor P18 Arabic Industry practitioner
P8 Indic languages Trust and safety vendor P19 Arabic Industry practitioner
P9 Quechua Linguist P20 Arabic Industry practitioner
P10 Quechua PhD student P21 Arabic PhD student
P11 Quechua Professor P22 Arabic Startup founder